OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) Let f(x) = x2 - 3x - 7. Find f(-3). 11 8 -7 -10 Question 2 (Multiple Choice Worth 1 points) Let f(x) = x2 - 16. Find f-1(x). 4square root of x ±square root of x plus 16 x squared over 16 1 over quantity of x squared minus 16 Question 3 (Multiple Choice Worth 1 points) Let f(x) = 2x - 6 Solve f-1(x) when x = 2. -5 4 -2 7 Question 4 (Multiple Choice Worth 1 points) Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (anonymous):

1.f(-3)=(-3)^2-3(-3)-7 =9+9-7 =11 2. 4 square root of x square root of x plus 16 x squared over 16 1 over quantity of x squared minus 16 3. Start with y = 2x - 6 Switch variables to get x = 2y - 6 Solve for y to get y = (1/2)(x+6) Your inverse function is f^-1(x) = (1/2)(x+6) and f^-1(2) = (1/2)(2+6) = 4 4. f(x) = 2x + 2 assuming you mean inverse: f-1(x) = (x - 2)/2 f-1(4) = (4-2)/2 = 1

OpenStudy (anonymous):

@dantheman26

OpenStudy (anonymous):

Hope I helped XD

OpenStudy (anonymous):

THANK YOU!!!!!